# GIM: A Million-scale Benchmark for Generative Image Manipulation Detection and Localization

### v1

@@ Abstract
The extraordinary ability of generative models emerges as a new trend in image editing and generating realistic images, posing a serious threat to the trustworthiness of multimedia data and driving the research of image manipulation detection and location(IMDL). However, the lack of a large-scale data foundation makes IMDL task unattainable. In this paper, a local manipulation pipeline is designed, incorporating the powerful SAM, ChatGPT and generative models. Upon this basis, We propose the GIM dataset, which has the following advantages: 1) Large scale, including over one million pairs of AI-manipulated images and real images. 2) Rich Image Content, encompassing a broad range of image classes 3) Diverse Generative Manipulation, manipulated images with state-of-the-art generators and various manipulation tasks. The aforementioned advantages allow for a more comprehensive evaluation of IMDL methods, extending their applicability to diverse images. We introduce two benchmark settings to evaluate the generalization capability and comprehensive performance of baseline methods. In addition, we propose a novel IMDL framework, termed GIMFormer, which consists of a ShadowTracer, Frequency-Spatial Block (FSB), and a Multi-window Anomalous Modelling (MWAM) Module. Extensive experiments on the GIM demonstrate that GIMFormer surpasses previous state-of-the-art works significantly on two different benchmarks. Code page: GIM.
Keywords:
Dataset and Benchmark Image Manipulation Detection and Location Forgery Traces Extraction Multi-scale Feature Learning
11footnotetext: * These authors contributed equally to this research, which was done during Yirui Chen and Quan Zhang’s internship at Huawei Noah’s Ark Lab. ‡ Project lead † Corresponding authors
1 Introduction
@@
Figure 1: Example image from the GIM dataset. Our dataset includes images manipulated by three state-of-the-art generators: Stable-Diffusion, GLIDE, and DDNM. The first column shows authentic images, while the second column presents forgery masks. The third column displays forged images.
@@
Images are one of the most essential media for information transmission in modern society, which is widely spread on public platforms such as news and social media. With the rapid advancement of generative methods [11, 52, 16, 4], such natural information can be more easily tampered with for specific purposes such as tampering with an object or person. The image of this class is particularly convincing since its visual comprehensibility, leads to serious information security risks in social areas. For instance, an AI-generated image of black smoke billowing from what appeared to be a government building near the Pentagon set off investor fears, sending stocks tumbling [46]. Therefore, it is of utmost urgency to develop effective methods to examine whether an image is modified by generative models, meanwhile identifying the exact location of the tampering. However, the traditional image manipulation detection and location(IMDL) dataset [55, 13, 60] overlook the powerful generative models and generative IMDL dataset are limited with scale, thus not sufficient to comprehensively evaluate the performance of IMDL methods and benefits the IMDL community.
@@
To this end, we propose a million-level generative-based IMDL dataset, termed GIM, to provide a reliable database for AI Generated Content (AIGC) security. GIM leverage the diffusion model [44, 52, 59] and SAM [32, 51], with ImageNet [10] and VOC [14] as the input. SAM is leveraged to locate the tampering region, and diffusion models paint the reasonable content in the tampering area. The excellent diffusion model ensures dataset fidelity and rich categories guarantee the diversity of images in GIM. GIM contains a total of over one million generative tampered images. Based on this, the IMDL methods are evaluated and benchmarked. To develop an appropriate benchmark scale, we explore the impact of different amounts of training manipulated data. The final benchmark contains about 300k manipulated images with their tampering masks. To simulate the real situation, we investigate the effect of degradation and apply the images to the random degradation method. Two benchmark settings are designed to evaluate the comprehensive and generalization. In addition, we recreated existing IMDL methods in a fair manner, providing the basis for future development. Overall, as a generative IMDL benchmark, GIM possesses the following advantages: 1) GIM has a large and reasonable data scale, including rich image categories and contents. 2) GIM contains various generative manipulation models and tasks. 3) GIM proposes two settings for verifying generalization capability and comprehensive performance in existing methods.
@@
The existing methods emphasize traditional image manipulations also known as “cheapfakes”. Generative manipulations introduce lethal alterations in content with no apparent frequency or structural inconsistency. To address the above issue, we introduce GIMFormer, a transformer-based framework for generative IMDL. The ShadowTracer is designed to embed the nuanced artifacts inherent in generative tampering and serves as prior information. The Frequency-Spatial Block harvests the manipulation clues in frequency and spatial domains. Furthermore, the Multi Windowed Anomalous Modelling Module captures local inconsistencies at different scales to refine the features. By doing so, our model extracts features from both the RGB and learned tampering traces map, captures details from the frequency domain and spatial domain, and models inconsistencies at different scales for precise manipulation detection and location.
We conduct experiments on our proposed GIM. Both the qualitative and quantitative results demonstrate that GIMFormer can significantly outperform previous state-of-the-art methods in terms of both comprehensive capability and generalization ability.
In summary, our main contributions are as follows:
  • We build a local manipulation pipeline and construct a comprehensive large-scale dataset with various SOTA generative models and manipulation tasks to further facilitate the research on the IMDL task.
  • We investigate the impact of data scales and degradation on IMDL tasks to select an appropriate scale and simulate real-world scenarios. Based on this, we construct a reasonable benchmark and two settings to evaluate the generalization and comprehensive capabilities of IMDL methods.
  • We propose GIMFormer, a generative IMDL framework consisting of ShadowTracer to embed subtle traces, Frequency-Spatial Block to extract forgery features in the frequency and spatial domains and Multi-Windowed Anomalous Modelling Module to capture pixel-level inconsistencies at multi scales.
@@
  • Extensive experiments show that GIMFormer achieves state-of-the-art IMDL performance in terms of generalization and comprehensive capabilities.
2 Related Work
2.1 Image Forensic Datasets
@@
The research community has dedicated significant efforts to establishing robust datasets for image forensics. Early datasets primarily focus on one type of manipulation. The Columbia dataset  [43, 24] is the pioneering dataset focusing on splicing forgeries. CASIA v1.0 and CASIA v2.0  [13] are the first to incorporate multiple types of manipulations within a single dataset, with forged images manually crafted using Adobe Photoshop. The Wild Web [66] dataset collects forged images from the Internet, significantly surpassing the scale of previous datasets. NIST [17] Datasets provide an extensive collection of datasets serving as a crucial standard for assessing media tampering detection methods. DEFACTO [42]is automatically generated from the Microsoft COCO dataset, encompassing four categories of forgeries: Copy and Move, Splicing, Object Removal, and Morphing. IMD2020 [45] provides a range of locally manipulated images generated through manual operations or random slicing, while also collecting online images without apparent manipulation traces. Recently, HiFi-IFDL [20] construct a hierarchical fine-grained dataset containing some representative forgery methods. AutoSplice [29] leverage the capabilities of large-scale language-image models like DALL-E2  [48] to facilitate automatic image editing and generation. CocoGlide [19] contains 512 images using the GLIDE diffusion model. Some datasets focused exclusively on facial manipulations [54, 18, 9] or entirely synthesized images[71, 56, 3].
@@
However, these datasets have certain limitations, such as small data sizes and limited manipulation techniques or tasks. Recent advances in generative models [44, 23, 52] have demonstrated remarkable abilities in generative-based manipulation. These developments have led to the emergence of open-source projects [51, 65, 15] for manipulation. Leveraging the power of these models, we introduce GIM, a large-scale generative-based manipulation dataset.
@@
Figure 2: An overview of the dataset generation. With the input of the original image and the desired region of interest, classification attribution or mouse input as a User query, the manipulation mask is extracted by the zero-shot segmentation model SAM. Besides, interacting with ChatGPT, we organize the tampering prompt by combining replacement classes. The final generations are produced by generative models with the input of the original image, tampering mask and tampering prompts.
2.2 Image Manipulation Detection and Localization
@@
Early studies on natural image manipulation localization mainly focus on detecting a specific type of manipulation [7, 27, 34, 50]. Due to the exact manipulation type is unknown in real-world scenarios, most state-of-the-art methods [40, 19, 58, 64, 38, 28, 61] primarily concentrate on general manipulation. RGB-N [70] propose a dual-stream network for manipulation detection. One stream focuses on extracting RGB features and the other leverages noise features to model the disparities. MVSS-Net [12] utilizes a two-stream CNN to extract noise features and employs the Dual Attention Module to merge the output of the two-stream CNN. PSCC-Net  [40] extracts hierarchical features with a top-down path and detects whether the input image has been manipulated using a bottom-up path. ObjectFormer [58] uses object prototypes to model object-level consistencies and find patch-level inconsistencies to detect the manipulation. However, these methods focus on "cheapfake" detection and encounter challenges when applied to generative manipulation due to the tampering traces are subtle and the lack of an understanding of generative manipulation patterns. Certain works [26, 6, 9, 39, 2] focus on localizing manipulations using generative models, yet these methods primarily concentrate on human faces.
@@
In this work, we leverage a deep network to embed the subtle artifacts inherent in generative tampering as a prior trace map. we then employ a dual network to fuse the trace map and RGB image, combining frequency and spatial information and capturing pixel-level inconsistencies at multiple scales to detect and locate the manipulation.
Table 1: Summary of previous image manipulation datasets and GIM. We showcase the number of data entries within each dataset, the image content, image sizes and the manipulation techniques they encompass. $\star$ denotes that HiFi-IFDL is composed of multiple datasets, including entirely synthesized images, traditionally manipulated images and fake face images.
@@
# Authetic Image # Forged Image Traditional Gen.Reprint Gen.Removal
FaceForensics++ [54] Face 480p-1080p 1000 4000
DFDC [9] Face 240p-2160p 19,154 100,000
DeeperForensics [30] Face 1080p 50,000 10,000
Columbia Gray [43] General 128 $\times$ 128 933 912
CASIA V2.0  [13] General 320 $\times$ 240-800 $\times$ 600 7,200 5,123
IMD2020 [45] General 1062 $\times$ 866 35,000 35,000
Coverage [60] General 400 $\times$ 486 100 100
AutoSplice  [29] General 256 $\times$ 256 - 4232 $\times$ 4232 3,621 2,273
HiFi-IFDL[20] General - - 1,000,000$\star$
CocoGlide [19] General 256$\times$256 - 512
@@
3 Dataset and Benchmark Construction
In Section 3.1, we propose an automatic data synthesis pipeline to produce generative manipulated images efficiently with unlabeled images. Leveraging this pipeline, we construct a comprehensive large-scale dataset. To build a reasonable scale benchmark for evaluating the IMDL methods, we delve into the impact of training data scale in Section 3.2. To emulate real-world conditions and establish an objective benchmark, the manipulated images and original authentic images are subjected to three random degradations (JPEG compression, downsampling, and Gaussian blur), as detailed in Section 3.3). GIM benchmark comprises over 320k manipulated images paired with authentic images (100 distinct labels in ImageNet for each generative model) for algorithm evaluation. The generated images are shown in Figure 1 with their original images and tampering masks. The specific composition of GIM can be found in the Supplementary Materials. In Section 3.4, we introduce the criteria and settings used to evaluate the performance of IMDL methods.
3.1 Generation and Details of the dataset
To construct a generative-based IMDL dataset, a large-scale natural dataset is urgently desired. ImageNet [10] and VOC [14] are awesome datasets, chosen as the starting point for our research due to the following advantages: 1) Large and diverse. 2) Wide category coverage. To sum up, we regard these datasets as the database for generative manipulation. In addition to possessing various scenarios, the proposed dataset utilizes a wealth of tampering methods. Specifically, we reprint the target class by generative inpainting [52, 44] method or remove the destination region using the model [59].Benefiting from open-source projects [33, 51], we develop our data generation pipeline.
@@ 3.1 Generation and Details of the dataset
Figure 2 illustrates the overall process. Firstly, with the classification attribute or User query, the local manipulation mask is extracted by the zero-shot segmentation network [33]. For reprint tampering, the image category is embedded in the replacement prompt and interacts with ChatGPT, which returns an approximate category. The approximate category is then embedded into the inpainting prompt. Combined with the original image, manipulation mask, and inpainting prompt, generative models generate the reprint generative tampered result. For removal tampering, only the original image and manipulation mask are required for the generated model. The GIM Dataset utilizes the entire ImageNet dataset and the Stable-Diffusion generator to construct a comprehensive dataset, providing a reliable database and laying the foundation for the research below.
3.2 Analysis of Benchmark Scale
@@
With the data generation pipeline described above, we are qualified to generate millions, tens of millions, or even billions of data. Nonetheless, blindly increasing the amount of data does not improve the algorithm performance, but may lead to data redundancy. Therefore, taking data generated by the Stable Diffusion as an example, we explore the influence of data volume and category volume on baseline classification [21] and segmentation algorithms [63]. Training sets are the different scales of the training subset, and the validation is the same test set. As shown in Table 2, the metrics of classification and segmentation are improved as the dataset scale increases. When the scale reaches 180K, the algorithm performance is almost saturated. No matter whether the image class or data scale is added, the metrics are stagnant. Experiments demonstrate that increasing the amount of data or category brings negligible benefits when the data tends to be saturated. According to the analysis above, for the training data, the GIM benchmark selects 100 different labels from ImageNet for each generation model to create tampered images. For the test data, the GIM benchmark uses the entire test dataset for each generation model.
Table 2: Results of classification and segmentation on different scale subsets. Classification uses Resnet-50 and segmentation uses SegFormer-b0.
Total Num. of Image Image Classes Image per Class Metrics
[…]
2,800 10 280 0.6306 0.2550 0.7950
28,000 100 280 0.7480 0.3247 0.7995
180,000 100 1800 0.9129 0.5289 0.8699
360,000 200 1800 0.9131 0.5291 0.8701
500,000 500 1000 0.9131 0.5291 0.8703
Table 3: Influence of different degradation method on classification (Resnet-50) and segmentation (SegFormer-b0) .
Degradation Metrics
@@
- 91.29 52.89
JPEG 83.10 45.74
Gaussian Blur 90.94 40.14
Downsample 88.34 37.12
3.3 Post-processing of Degradation Method
@@
After being spread on the Internet, images will encounter various post-processings, such as JPEG compression, downsampling, Gaussian blur, etc. These transformations can pose challenges for image forensics methods. Figure 3.2 investigates the impact of degradation. The baseline models are trained on the clean data generated by the Stable-Diffusion training set and tested on the same test set with a specific degradation method. Experimental results denote that these degradation methods make identification more difficult. To shorten the distance from reality and build an objective benchmark, the random degradation method (JPEG compression, downsampling, Gaussian blur) is performed on the dataset. The details of the parameters can be found in Supplementary Materials.
3.4 Benchmark Settings
Dataset: The datasets used for the research of generative IMDL tasks are lacking. This makes direct comparisons and accurate measurement difficult. We use three cross-generator datasets GIM-SD (data from ImageNet manipulation by Stable Diffusion, similarly hereinafter), GIM-GLIDE, GIM-DDNM and a cross-distribution dataset GIM-VOC(data from VOC manipulation by Stable Diffusion) to present benchmarking results.
@@ 3.4 Benchmark Settings
Metrics: We evaluate the performance of the proposed method on both the image manipulation detection task and localization task. For classification results, we use Accuracy (Cls.acc) as our evaluation metric, while for localization, the pixel-level AUC and F1 score on manipulation masks are adopted. Since binary masks and detection scores are required to compute F1 scores, we adopt the Equal Error Rate (EER) threshold to binarize them.
@@
Settings: Two settings are proposed to verify the comprehensive and generalization performance of the algorithms. In the cross-generator generalization setting, the models are trained on the GIM-SD training set and tested on the GIM-GLIDE, GIM-DDNM and GIM-VOC test set to explore the generalization IMDL performance, as shown in Table 5. In the mix-generator comprehensive setting, the models are jointly trained on the GIM-SD, GIM-GLIDE and GIM-DDNM training set and tested on the correspondence test dataset respectively to evaluate the comprehensive IMDL performance, as shown in Table 4.
4 Method
[…]
@@ 4 Method
Figure 3: GIMFormer architecture. The ShadowTracer takes the RGB image $x$ to extract the learned manipulation trace map $t$. The encoder uses both the RGB input and trace map $t$ to extract pyramidal features $F_{i}$ ($i\in[1,4]$) across four stages. The four-stage fused features $F_{i}$ are forwarded to the decoder for generative manipulation detection and location.
4.1 ShadowTracer
@@
Prior manipulation detection methods mainly focus on ”cheapfake” scenes and rely on visible traces. These artifacts include distortions and sudden changes caused by manipulation of the image structure. However, generative tampering makes significant alterations to the content with no apparent frequency or structural inconsistency. As shown in Figure 4, these subtle traces are displayed with inherent patterns, not visible traces with inconsistent edges.
@@
ShadowTracer aims to capture the inherent characteristics and subtle traces of the generative models. For a doctored image, our objective is to learn a mapping $g_{\phi}$ to map the tampered image to its latent disturbed pixel values, where $g_{\phi}$ represents a neural network with trainable parameters $\phi$. Our key observation is that the differences introduced by generative models in data distribution exhibit inherent patterns, and deep neural networks can attempt to reconstruct these variations. At the training stage, we generate pairs of the input image $x_{i}$ and the generative tampered image $G(x_{i})$, the manipulation trace can be calculated by $t_{i}=G(x_{i})-x_{i}$. The objective function for training $g_{\phi}$ can be formulated as
$\min_{\bm{\phi}}\Big{\{}\mathcal{L}_{r}(g_{\bm{\phi}}\left(G\left(\mathbf{x}_{i}\right))\right),t_{i})\Big{\}}$ (1)
where $\mathcal{L}_{r}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$.
@@
Furthermore, the mapping network should satisfy two properties: detecting subtle tampering traces and being robust to various real-world image degradations. For this reason, image pairs are generated by mixing original and manipulated images and incorporating diverse degradation operations at the training stage. Specifically, given an input image $I$, we segment a portion and perform generative manipulation to obtain $I_{m}$. The alpha blendering [47] is utilized to the original and manipulated images to obscure obvious tampering traces. Following this, we subject the images to degradation operations such as blurring, and JPEG compression to obtain the final manipulated image. The network is trained on 64$\times$64 pixel patches randomly sampled from the dataset, adapting the loss in eq. 1.
4.2 Frequency-Spatial Block

### v2

@@ Abstract
The extraordinary ability of generative models emerges as a new trend in image editing and generating realistic images, posing a serious threat to the trustworthiness of multimedia data and driving the research of image manipulation detection and location (IMDL). However, the lack of a large-scale data foundation makes the IMDL task unattainable. In this paper, we build a local manipulation data generation pipeline that integrates the powerful capabilities of SAM, LLM, and generative models. Upon this basis, we propose the GIM dataset, which has the following advantages: 1) Large scale, GIM includes over one million pairs of AI-manipulated images and real images. 2) Rich image content, GIM encompasses a broad range of image classes. 3) Diverse generative manipulation, the images are manipulated images with state-of-the-art generators and various manipulation tasks. The aforementioned advantages allow for a more comprehensive evaluation of IMDL methods, extending their applicability to diverse images. We introduce the GIM benchmark with two settings to evaluate existing IMDL methods. In addition, we propose a novel IMDL framework, termed GIMFormer, which consists of a ShadowTracer, Frequency-Spatial block (FSB), and a Multi-Window Anomalous Modeling (MWAM) module. Extensive experiments on the GIM demonstrate that GIMFormer surpasses the previous state-of-the-art approach on two different benchmarks.
Introduction
@@
Figure 1: Example images from the GIM dataset. Our dataset includes images manipulated by three state-of-the-art generators: Stable-Diffusion, GLIDE, and DDNM. Three columns display authentic images, manipulation masks and manipulated images.
@@
Images are one of the most essential media for information transmission in modern society and they are widely spread on public platforms such as news and social media. With the rapid advancement of generative methods (Dhariwal and Nichol 2021; Rombach et al. 2022), such natural information can be more easily manipulated for specific purposes such as tampering with an object or person. The image of this class is particularly convincing since its visual comprehensibility, leading to serious information security risks in social areas. Therefore, it is of utmost urgency to develop methods to detect whether an image is modified by generative models and identify the exact location of the manipulation. However, traditional image manipulation detection and location (IMDL) datasets (Dong, Wang, and Tan 2013; Wen et al. 2016) overlook the powerful generative models and generative IMDL (Jia et al. 2023; Guillaro et al. 2023) datasets are limited with scale, it is thus not sufficient to comprehensively evaluate the performance of IMDL methods and benefit the IMDL community.
@@
To this end, we propose a million-level generative-based IMDL dataset, termed GIM dataset, to provide a reliable database for AI Generated Content (AIGC) security. GIM leverages the generative models (Ho, Jain, and Abbeel 2020; Yin et al. 2025) and SAM (Kirillov et al. 2023a), with the images in ImageNet (Deng et al. 2009) and VOC (Everingham et al. 2010) as the input. SAM is utilized to locate the tampering region, and generative models paint the reasonable content in the tampering area. GIM contains a total of over one million generative manipulated images. To develop an appropriate benchmark scale, we explore the impact of different amounts of training manipulated data. The final benchmark contains about 320k manipulated images with their tampering masks for training and testing. To simulate the real situation, we investigate the effect of degradation and apply random degradation to these benchmark images. Based on this GIM benchmark, the IMDL methods are evaluated and benchmarked. Overall, GIM possesses the following advantages: 1) GIM has a large and reasonable data scale, including rich image categories and contents. 2) GIM contains various generative manipulation models and tasks. 3) GIM proposes two settings for verifying the performance and generalization ability of IMDL methods.
@@
Existing methods emphasize traditional image manipulations also known as “cheapfakes”. However, generative manipulations introduce lethal alterations in content with no apparent frequency or structural inconsistency. To address the above issue, we introduce GIMFormer, a transformer-based framework for generative IMDL. The ShadowTracer is designed to embed the nuanced artifacts inherent in generative tampering and serves as the prior information. The Frequency-Spatial Block captures the manipulation clues in the frequency and spatial domains. Furthermore, the Multi Windowed Anomalous Modelling module captures local inconsistencies at different scales to refine the features. GIMFormer extracts features from both the RGB and learned tampering trace maps, capturing details from both the frequency and spatial domains while modeling inconsistencies at different scales for precise manipulation detection and localization. We conduct experiments on our proposed GIM. Both the qualitative and quantitative results demonstrate that GIMFormer can outperform previous state-of-the-art methods. In summary, our main contributions are as follows:
  • We build a data generation pipeline and construct a large-scale dataset for IMDL tasks.
  • We investigate the impact of data scales and degradation, constructing a comprehensive benchmark with two evaluation settings for IMDL methods evaluation.
@@
  • We propose a framework named GIMFormer for generative IMDL task. Extensive experiments demonstrate that GIMFormer achieves state-of-the-art performance.
Related Work
Image Forensic Datasets
@@
Early datasets (Ng, Chang, and Sun 2004) primarily focus on one type of manipulation. CASIA  (Dong, Wang, and Tan 2013) first incorporates multiple types of manipulations with forged images manually crafted using Adobe Photoshop. The Wild Web dataset (Zampoglou, Papadopoulos, and Kompatsiaris 2015) collects forged images from the Internet, surpassing previous datasets in the scale. NIST (Guan et al. 2019) dataset provides an extensive collection of datasets serving as a crucial standard for assessing media tampering detection methods. IMD2020 (Novozamsky, Mahdian, and Saic 2020) provides locally manipulated images generated through manual operations or random slicing and online images without obvious manipulation traces. Recently, HiFi-IFDL (Guo et al. 2023) constructs a hierarchical fine-grained dataset containing some representative forgery methods. AutoSplice (Jia et al. 2023) leverages large-scale language-image models like DALL-E2  (Ramesh et al. 2021) to facilitate automatic image editing and generation. CocoGlide (Guillaro et al. 2023) contains images manipulated by the diffusion model. Besides, other datasets focus on facial manipulations (Rossler et al. 2019; Guarnera et al. 2022) or entirely synthesized images(Zhu et al. 2023; Yan et al. 2024; Verdoliva and Cozzolino 2022).
@@
The datasets mentioned above have limitations such as small data sizes and limited manipulation techniques. Recent advances in generative models have demonstrated remarkable manipulation abilities. Leveraging these models, we introduce GIM, a large-scale dataset that incorporates various recent generative manipulation techniques.
@@
Figure 2: An overview of the dataset generation. Given the original image and a user query (classification attribution or mouse input), the manipulation mask is extracted using SAM. Tampering prompts are then organized with LLM by combining replacement classes. The final generations are produced by generative models with the image, tampering mask and prompts.
Image Manipulation Detection and Localization
@@
Early studies on natural image manipulation localization mainly focus on detecting a specific type of manipulation (Cozzolino, Poggi, and Verdoliva 2015; Yin et al. 2023). Due to the exact manipulation type is unknown in real-world scenarios, most state-of-the-art methods (Liu et al. 2022; Ying et al. 2023) primarily concentrate on general manipulation. MVSS-Net (Dong et al. 2022) uses a two-stream CNN to extract noise features and employs dual attention to fuse their outputs. PSCC-Net (Liu et al. 2022) extracts hierarchical features in a top-down manner and it detects manipulations in a bottom-up manner. Motivated by the power of transformer(Zhang et al. 2024; Zhang and Qi 2024; Lu et al. 2024a, b; Chen et al. 2023), ObjectFormer (Wang et al. 2022) uses object prototypes to model object-level consistencies and find patch-level inconsistencies to detect the manipulation. However, these methods focus on “cheapfake” detection and encounter challenges when applied to generative manipulation. Certain works (Huang et al. 2022; Chai et al. 2020; Liu et al. 2024) localize manipulations using generative models but mainly focus on human faces.
@@
In this work, we focus on the generative IMDL task. We leverage a deep network to embed the subtle artifacts inherent in generative tampering as a prior trace map. A dual network fuses the trace map and RGB image, combining frequency and spatial information and capturing pixel-level inconsistencies to detect and locate the manipulation.
Dataset and Benchmark Construction
In this section, We propose an automatic pipeline to generate manipulated images from unannotated data. Leveraging this pipeline, we construct the comprehensive large-scale GIM dataset. To build a reasonable benchmark, we conduct preliminary experiments focusing on two key aspects: data scale and image degradation. First, we investigate the impact of the training data scale to determine an appropriate size for the GIM benchmark. Second, to reveal real-world scenarios, the manipulated and original images are subjected to three random degradations. Finally, the GIM benchmark comprises over 320k manipulated images paired with authentic images for algorithm training and evaluation. The example images are shown in Figure 1 with their original images and tampering masks. Finally, we introduce the criteria and settings used to evaluate the IMDL methods.
@@
# Authentic Image # Forged Image Traditional Gen.Reprint Gen.Removal
FaceForensics++ (Rossler et al. 2019) Face 480p-1080p 1,000 4,000
DFDC (Dang et al. 2020) Face 240p-2160p 19,154 100,000
DeeperForensics (Jiang et al. 2020) Face 1080p 50,000 10,000
Columbia Gray (Ng, Chang, and Sun 2004) General 128 $\times$ 128 933 912
CASIA V2.0  (Dong, Wang, and Tan 2013) General 320 $\times$ 240-800 $\times$ 600 7,200 5,123
IMD2020 (Novozamsky, Mahdian, and Saic 2020) General 1062 $\times$ 866 35,000 35,000
Coverage (Wen et al. 2016) General 400 $\times$ 486 100 100
AutoSplice  (Jia et al. 2023) General 256 $\times$ 256 - 4232 $\times$ 4232 3,621 2,273
HiFi-IFDL (Guo et al. 2023) General - - 1,000,000 (200,000)
CocoGlide (Guillaro et al. 2023) General 256$\times$256 - 512
@@
Table 1: Summary of image manipulation datasets. Image content, image sizes, and manipulation techniques are reported. HIFI-IFDL includes mainly entirely synthesized images with a small portion of traditional manipulation or face image manipulation.
Data Generation Pipeline
@@
Benefiting from open-source projects (Kirillov et al. 2023b; Ren et al. 2024), we develop our data generation pipeline. Figure 2 illustrates the overall process. Our pipeline includes two types of manipulation: reprinting the target class using a generative inpainting method Stable Diffusion (Rombach et al. 2022) and GLIDE  (Nichol et al. 2021) or removing the destination region using a specific removal method DDNM (Wang, Yu, and Zhang 2022). Given an arbitrary image, the process begins by extracting the local manipulation mask using a zero-shot segmentation network SAM, guided by either classification attributes or user queries. For reprint tampering, the image category is embedded in the replacement prompt and interacts with LLM (Achiam et al. 2023), which returns an approximate category. The approximate category is then embedded into the inpainting prompt. By combing with the original image, manipulation mask and inpainting prompt, generative models generate the reprint tampered result. For removal tampering, only the original image and manipulation mask are required for the model.
The GIM dataset utilizes images from the ImageNet dataset  (Deng et al. 2009) and VOC dataset  (Everingham et al. 2010). These two datasets contain a large number of diverse images with wide category coverage, providing a reliable database and laying the foundation for future research.
Analysis of the Benchmark Scale
@@
With the data generation pipeline, we are qualified to generate large amounts of manipulated images. Nonetheless, blindly increasing the amount of data does not improve the algorithm performance, while it may lead to data redundancy. Therefore, taking data generated by the Stable Diffusion as an example, we explore the influence of data volume and category volume on baseline classification (He et al. 2016) and segmentation algorithms (Xie et al. 2021). Training sets vary in the scale, while the validation uses the same test set. As shown in Table 2, the metrics of classification and segmentation are improved as the dataset scale increases. When the scale reaches 180K, the algorithm performance almost saturates. All the metrics remain almost unchanged with either the image class number or the data scale increased. Experiments demonstrate that increasing the amount of data or category brings negligible benefits when the data tends to be saturated. According to the analysis, the GIM benchmark uses 100 labels from ImageNet to generate tampered images for training and employs all the test sets from ImageNet and VOC for evaluation.
Total Num. of Image Image Classes Image per Class Metrics(%)
Cls.Acc Seg.F1 Seg.AUC
2,800 10 280 63.1 25.5 79.5
28,000 100 280 74.8 32.5 79.9
180,000 100 1800 91.3 52.9 86.9
360,000 200 1800 91.3 52.9 87.0
500,000 500 1000 91.3 52.9 87.0
Table 2: Dataset Scale Experiment: Effect of Dataset Scale on the Performance of Base Models (SegFormer-b0, ResNet-50).
Degradation Metrics(%)
@@
- 91.3 52.9
JPEG compression 83.1 45.7
Gaussian blur 90.9 40.1
Downsample 88.3 37.1
Table 3: Degradation Experiment: Impact of Degradation on Base Models (SegFormer-b0, ResNet-50).
Post-processing of Degradation Method
@@
When uploaded to the Internet, images will encounter various post-processing. These transformations can pose challenges for image forensics methods. Table 3 investigates the impact of degradation. The baseline models are trained on the clean data and tested on the test set with a single degradation. Experimental results show that these degradations make the identification difficult. To reveal the real-world scenarios, random degradation (JPEG compression, downsampling and Gaussian blur) is performed on the dataset.
Benchmark Settings
Benchmark Description: GIM consists of four subsets. GIM-SD, GIM-GLIDE and GIM-DDNM which contain data from ImageNet manipulated by Stable Diffusion, GLIDE and DDNM, respectively. Additionally, there is a cross-distribution subset GIM-VOC which contains data from VOC manipulated by Stable Diffusion.
@@
Metrics: We evaluate the performance of the IMDL method on both the image manipulation detection task and localization task. For the detection task, we use accuracy (Cls.acc) as our evaluation metric. For the localization task, the pixel-level AUC and F1 score are adopted.
@@
Settings: Two settings are proposed to evaluate the performance and generalization. In the mix-generator setting, the models are jointly trained on the GIM-SD, GIM-GLIDE and GIM-DDNM training set and tested on the correspondence test dataset respectively to evaluate the performance. In the cross-generator setting, the models are trained on the GIM-SD training set and tested on the GIM-GLIDE, GIM-DDNM and GIM-VOC test sets to explore the generalization.
Method
To address the challenges of generative manipulation, we propose GIMFormer that adopts a dual encoder and decoder architecture. Our framework includes several components: the ShadowTracer, the Frequency-Spatial Block (FSB), and the Multi Windowed Anomalous Modeling (MWAM) module. Figure 3 gives an overview of the framework. For the RGB image input $x$, we first extract its learned trace map $t$. Then, both $x$ and $t$ are fed into a two-branch network, where the four-stage structure is used to extract pyramid features $F_{i}$ ($i\in[1,4]$). The RGB branch is composed of FSB, Transformer Block (Xie et al. 2021) and WMAM. The tracer branch consists of a Transformer Block and WMAM. In the fusion step, the feature rectification module (FRM) and feature fusion module (FFM) (Zhang et al. 2023) are used for feature fusion. The four-stage fused features are forwarded to the decoder for final detection $\hat{y}$ and location $\hat{M}$.
@@
Figure 3: GIMFormer architecture. ShadowTracer extracts trace map $t$ from the input image $x$. The encoder combines $x$ and $t$ to generate pyramid features $F_{i}$ across four stages, which are sent to the decoder for manipulation detection and localization.
ShadowTracer
@@
Prior manipulation detection methods mainly focus on “cheapfake” and rely on visible traces. These artifacts include distortions and sudden changes caused by manipulation of the image structure. However, generative tampering makes significant alterations to the content with no apparent frequency or structural inconsistency. As shown in Figure 4, these subtle traces are displayed with inherent patterns that are not visible traces with inconsistent edges.
@@
ShadowTracer aims to capture the inherent characteristics and subtle traces of the generative models. For a manipulated image, our objective is to learn a mapping $g_{\phi}$ to map the tampered image to its latent disturbed pixel values, where $g_{\phi}$ represents a neural network with trainable parameters $\phi$. Our key observation is that the differences introduced by generative models in data distribution exhibit inherent patterns, and deep neural networks can attempt to reconstruct these variations. At the training stage, we generate pairs of the image $x_{i}$ and the tampered image $G(x_{i})$, the manipulation trace can be calculated by $t_{i}=G(x_{i})-x_{i}$. The objective function for training $g_{\phi}$ can be formulated as :
$\min_{\boldsymbol{\phi}}\Big{\{}\mathcal{L}_{r}(g_{\boldsymbol{\phi}}\left(G\left(\mathbf{x}_{i}\right))\right),t_{i})\Big{\}}$ (1)
@@
where $\mathcal{L}_{r}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$. Furthermore, the mapping network needs to detect subtle tampering traces and be robust to various real-world image degradations. For this reason, image pairs are generated by mixing original and manipulated images and incorporating diverse degradation operations at the training stage. Specifically, given an input image $I$, we segment the region of interest and perform generative manipulation to obtain $I_{m}$. The mix-up (Zhang et al. 2017a) strategy is utilized to the original and manipulated images to obscure obvious tampering traces. Following this, we subject the images to the degradations mentioned above to obtain the final manipulated image. The network is trained on 64$\times$64 pixel patches randomly sampled from the dataset and the loss in Eq. 1 is adopted.
Frequency-Spatial Block